The length of a garden is 20 yards more than its width. The area is 300 square yards. What are the dimensions? This is on a worksheet but my teacher didn't teach us how to do this. Help?
Answer by rothauserc(4718)

the area of a rectangle is length times width,
let x be the width of the rectangle and the length is x+20
now we can write the formula for the area as
300 = x * (x+20)
300 = x^2 +20x
subtract 300 from both sides of the =
x^2 +20x -300 = 0
this quadratic can be factored into
(x+30) * (x-10) = 0
therefore the two possible values for x are -30 and 10
we are looking for a positive value in this case, soooooooo
x = 10 and
the width is 10 yards and the length is 30 yards